OpenStudy (anonymous):

What is the solution set of the following system of equations? 5x - 3y - 10 = 0 x + y - 7 = 0 {(25/8, 31/8)}?

OpenStudy (anonymous):

@gypsy1274

OpenStudy (anonymous):

That is not what I am getting. Did you plug those values in to one of the equations to check?

OpenStudy (anonymous):

oh im sorry i have them switched

OpenStudy (anonymous):

Still not what I am getting.....

OpenStudy (anonymous):

Don't mind the idiot...(That would be me.)

OpenStudy (anonymous):

well here is preety much what i did first to fix out what either x or y equals. I chose to figure out what 7 equals using the second equation. Like this: x + y - 7 = 0 x + y = 7 y = 7 - x You can then plug this into the first equation so that you have only xs and numerical values and can find out what the value of x is. Like this: 5x - 3(7-x) - 10 = 0 You can do this because y = 7 - x as shown above 5x - 21 + 3x - 10 = 0 8x - 31 = 0 8x = 31 x = 31/8 Now that you know the numerical value of x you can plug it into the second equation to find the numerical value of y. Like this: 31/8 + y - 7 = 0 31/8 + y - (56/8) = 0 y - (25/8) = 0 y = 25/8 The solution is therefore x = 31/8 and y = 25/8.

OpenStudy (anonymous):

OK, I see my mistake. Anytime you want to check your answers on this type of question, just plug them in to one of the equations. That will tell you if you have the correct answer.

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

These answers are correct. \((\frac{31}{8},\frac{25}{8})\)

OpenStudy (anonymous):

ok so i was right the secound time

OpenStudy (anonymous):

Yes.